Obtain expressions in component form for the position vectors having the following polar coordinates.
(a) 12.2 m, 170° counterclockwise from the +x axis
R with arrow =
m

(b) 3.60 cm, 50.0° counterclockwise from the +x axis
R with arrow =
cm

(c) 20.0 in., 224° counterclockwise from the +x axis
R with arrow =
in.
1
Expert's answer
2020-04-13T10:17:02-0400

a)


"\vec R=(12.2\cos{170},12.2\sin{170})=(-12.0,2.12)\ m"

b)


"\vec R=(3.6\cos{50},3.6\sin{50})=(2.31,2.76)\ cm"

c)


"\vec R=(20\cos{224},20\sin{224})=(-14.4,-13.9)\ in"